OpenStudy (anonymous):

It is estimated that t years from now, the population of a certain suburban community will be p(t)=50 - 7/(2t+1) thousand people. At what rate, in people/year will the population be growing 3 years from now?

OpenStudy (anonymous):

dp/dt=14(2t+1)^-2 dp/dt=14(7)^-2 dp/dt=2/7 people/year

OpenStudy (anonymous):

I would have thought the first step would be 50-7(2t+t)^-1 then 43(2t+t)^-1 then -43(2t+t)^-2(2+1) then I get -43(3) (2(3)+(3))^-2.